OpenStudy (anonymous):

given the curve defined by x^2-siny=y+4 -Find d^2y/dx^2 in terms of x and y only

OpenStudy (anonymous):

taking derivative on both sides of equcation 2x-cos(y)y'=y' 2x=y'+cos(y)y' 2x=y'(1+cosy) 2x/(1+cosy)=y' at y=0 2x/(1+cos0)=y' 2x/(2)=y' x=y'